Calculate the complex power of the equivalent impedance of the following circuit, if the value of the current source A has a value of 2.3 A. Write the value of the apparent power in the solution, the angle will be verified in the procedure.

A cos(25t +30°) 40 800mH 4mF
Answer #1

The equivalent impedance of the following circuit can be given as-

1/Z = 1/R + 1/(jwL) + 1/(-j/wC)

=> 1/Z = 1/40 - j/(25 x 800mH) + j(25 x 4mF)

=> 1/Z = 0.025 - j0.05 + j0.1

=> 1/Z = 0.025 + j0.05

=> Z = (0.025 - j0.05)/(0.0252 + 0.052)

=> Z = 8 - j16

Now the complex power can be given as

S = Imax2/2Z

=> S = 2.32/2(8-j16)

=> S = 0.066125 + j0.13225 VA